OpenStudy (anonymous):

The surface area of a rectangular prism is 150 cm². If all the sides are tripled, what will the new surface area be?

OpenStudy (anonymous):

Area of rectangle = length * breadth

OpenStudy (anonymous):

So, length * breadth =150

OpenStudy (anonymous):

but if all the sides are tripled... what do i do?

OpenStudy (anonymous):

Now, find New Area = 3*length * 3*breadth=9*length*breadth=9*150=

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

oh yes thank you

OpenStudy (lgbasallote):

uhh...surface area...not area...different formulas...

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

Oh..... its a surface Ara

OpenStudy (anonymous):

isnt it Surface Area = 2Lw + 2wh + 2Lh = 150cm²

OpenStudy (anonymous):

Area

OpenStudy (lgbasallote):

that's why people shouldn't race in replying

OpenStudy (anonymous):

But same concept

OpenStudy (lgbasallote):

go on..redeem yourself @sauravshakya

OpenStudy (anonymous):

im going else where to find the answer. im noto going ot be used to sho whow smart people are. i just needed help.

OpenStudy (anonymous):

2Lw + 2wh + 2Lh = 150cm² 2(lw+wh+lh)=150 iw+wh+lh=75

OpenStudy (anonymous):

1. 2 wh + 2 wl + 2 lh = 150 2(wh + wl + lh) = 150 ===> (wh + wl + lh) = 75 triple: 2(9 wh +9 wl + 9lh) = 18 (wh + wl + lh) = 18*75 = 1,350

OpenStudy (anonymous):

Yep